OpenStudy (anonymous):

Can anyone differentiate this : y=Ln(tg(x/2))

hartnn (hartnn):

tg is tangent function ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

do you know chain rule ?

OpenStudy (anonymous):

I know how to do this type of problems but I can't get the right answer and I don't know where I go wrong

hartnn (hartnn):

so sho me your attempt, maybe i will find the error in your work...

hartnn (hartnn):

you know derivates of ln x, tg x and x/2 , right ?

OpenStudy (anonymous):

(1/(2cos^2 x/2))/tg x/2

hartnn (hartnn):

did u mean, \(\huge \dfrac{\dfrac{1}{2\cos^2(x/2)}}{\tan(x/2)}\) if yes, thats correct, just try to simplify a bit.

OpenStudy (anonymous):

Yes that's what I meant. And when I simplify it I get 1/(2sin x/2 * cos x/2)

OpenStudy (anonymous):

I don't know what to do from there on. None of the answers match.

hartnn (hartnn):

which is still correct now use sin 2x = 2 sin x cos x so, 2sin x/2 * cos x/2 will be just sin x and 1/sin x = cosec x is cosec x in choices ?

hartnn (hartnn):

and did u get what i did ?

OpenStudy (anonymous):

Oh yes, you are right! I forgot to use that equation at the end. The answer is 1/sinx. Thanks!! :)

hartnn (hartnn):

welcome ^_^